OpenStudy (anonymous):

x^2+y^2+2x-2y=-1 find vertex and radius

OpenStudy (anonymous):

this is an equation of a circle general form is (x-a)^2 + (y-b)^2 = r^2 where a,b is the centre and r = radius using completeing the square will let you figure these values out

myininaya (myininaya):

\[x^2+2x+y^2-2y=-1\] reorder \[x^2+2x+(2/2)^2+y^2-2y+(-2/2)^2=-1+(2/2)^2+(-2/2)^2\] whatever you add to one side you add to the other \[(x+1)^2+(y-1)^2=-1+1+1\] \[(x+1)^2+(y-1)^2=1\]

myininaya (myininaya):

center is (-1,1) radius is 1